OpenStudy (anonymous):

solve log5x-log13=0 round nearest hundredth if nec.

OpenStudy (anonymous):

\[\log \frac{ 5x }{13 }=0\] \[\frac{ 5x }{13 }=e ^{0}=1,x=?\]

OpenStudy (anonymous):

0.38 ?

OpenStudy (anonymous):

X=13/5=2.6

OpenStudy (anonymous):

so its 2.6 ?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

yw